# BioMedGPT-Mol: Multi-task Learning for Molecular Understanding and Generation

### v1

@@ Abstract
Molecules play a crucial role in biomedical research and discovery, particularly in the field of small molecule drug development. Given the rapid advancements in large language models, especially the recent emergence of reasoning models, it is natural to explore how a general-purpose language model can be efficiently adapted for molecular science applications. In this work, we introduce BioMedGPT-Mol, a molecular language model designed to support molecular understanding and generation tasks. By curating and unifying existing public instruction datasets, we have assembled a large-scale, comprehensive, and high-quality training dataset. The model is then fine-tuned through a meticulously designed multi-task learning framework. On a consolidated benchmark derived from LlaSMol, TOMG-Bench, and MuMOInstruct, BioMedGPT-Mol achieves remarkable performance. Our experimental results demonstrate that a general-purpose reasoning model can be effectively and efficiently post-trained into a professional molecular language model through a well-structured multi-task curriculum. Leveraging the power of it, we further explore retrosynthetic planning task, and the performance on RetroBench demonstrates its competitive capability of acting as an end-to-end retrosynthetic planner. We anticipate that our approach can be extended to other biomedical scientific domains.
@@ 1 Introduction
Retrosynthetic planning is a fundamental problem in organic chemistry, aiming to identify a set of commercially available starting materials and a sequence of reactions that can synthesize a given target molecule. To the best of our knowledge, LLMs alone have not previously been deployed as effective multi-step retrosynthetic planners. In this work, we further explore this challenging task with BioMedGPT-Mol using a three-stage SFT strategy and demonstrate its competitive capability of acting as a fully end-to-end retrosynthetic planner on RetroBench [liu2023fusionretro].
@@
  • We are the first to explore the retrosynthetic planning task using an LLM alone, and we introduce a three-stage SFT strategy that enables BioMedGPT-Mol to function as an end-to-end retrosynthetic planner.
@@ 7 Further Exploring Retrosynthetic Planning with BioMedGPT-Mol
In contrast to these representation-focused approaches, our work investigates the potential of fine-tuning LLMs as standalone generative planners via explicit Chain-of-Thought (CoT) reasoning. To the best of our knowledge, LLMs alone have not been deployed as successful multi-step retrosynthetic planners thus far. However, we argue that BioMedGPT-Mol, empowered by its deep molecular expertise and inherent context-learning capabilities, is capable of being fine-tuned as a standalone retrosynthetic planner, bypassing the need for any specialized search algorithms.
In this section, we validate the model’s potential by fine-tuning BioMedGPT-Mol on RetroBench dataset[liu2023fusionretro] to function as an end-to-end planner.
@@ 7.2 Performance on RetroBench
Figure 4: Summary of retrosynthetic planning results for exact match accuracy (%) on RetroBench.
A two-dimensional test-time scaling strategy is adopted, involving reasoning sampling and weighted beam search. We first employ repeated sampling with a non-zero temperature to generate $k_{s}$ diverse CoT trajectories. For each trajectory, we subsequently perform a partial beam search with a width of $k_{b}$ to decode the most probable reactant SMILES. The final predictions are ranked by aggregating the local beam scores across all sampled paths using the reciprocal rank sum:
$Score(y)=\sum_{j\in\mathcal{J}_{y}}\frac{1}{r_{j}}$ (2)
where $\mathcal{J}_{y}$ represents the set of occurrences for reactant $y$, and $r_{j}$ denotes its rank within the $j$-th beam. Figure 4 presents the exact-match accuracy ($\%$) across different methods, illustrating that our approach is highly competitive as a fully end-to-end retrosynthetic planner. Unlike NeuralSym [neuralsym], our method is template-free. RetroInText [kang2025retrointext] achieves an accuracy of $39.8\%$ by leveraging GPT-4, whereas our 8B-parameter model attains a comparable accuracy of $39.1\%$. We speculate that scaling the model size will yield substantial performance gains, which we leave for future investigation.
@@ 8 Conclusion
In this work, we introduce BioMedGPT-Mol, a molecular language model designed for molecular understanding and generation. To create a molecule-centric research assistant and harness the power of open-source reasoning models with rapid iterative upgrades, we design and implement a multi-task learning strategy for model fine-tuning. Our experimental results demonstrate its remarkable performance, showing that a general-purpose reasoning model can be effectively and efficiently post-trained into a specialized molecular language model through a well-structured multi-task curriculum. On the basis of that, we further explore the challenging retrosynthetic planning task. The proposed three-stage SFT strategy enables it to function as an end-to-end retrosynthetic planner.

### v2

@@ Abstract
Molecules play a crucial role in biomedical research and discovery, particularly in the field of small molecule drug development. Given the rapid advancements in large language models, especially the recent emergence of reasoning models, it is natural to explore how a general-purpose language model can be efficiently adapted for molecular science applications. In this work, we introduce BioMedGPT-Mol, a molecular language model designed to support molecular understanding and generation tasks. By curating and unifying existing public instruction datasets, we have assembled a large-scale, comprehensive, and high-quality training dataset. The model is then fine-tuned through a meticulously designed multi-task learning framework. On a consolidated benchmark derived from LlaSMol, TOMG-Bench, and MuMOInstruct, BioMedGPT-Mol achieves remarkable performance. Our experimental results demonstrate that a general-purpose reasoning model can be effectively and efficiently post-trained into a professional molecular language model through a well-structured multi-task curriculum. Leveraging these capabilities, we further apply the model to multi-step retrosynthetic planning, achieving state-of-the-art performance on RetroBench and demonstrating its superior efficacy as an end-to-end retrosynthetic planner. We anticipate that our approach can be extended to other biomedical scientific domains.
@@ 1 Introduction
Beyond the general capabilities covered by the multi-task framework, we specifically target retrosynthetic planning, a fundamental problem in organic chemistry aiming to identify a set of commercially available starting materials and a sequence of reactions that can synthesize a given target molecule. To the best of our knowledge, LLMs alone have not previously been deployed as effective multi-step retrosynthetic planners. In this work, we further advance BioMedGPT-Mol to tackle this challenge using a comprehensive training pipeline integrating SFT and RL. Consequently, our model achieves state-of-the-art performance on RetroBench [liu2023fusionretro], demonstrating its superior capability of acting as a fully end-to-end retrosynthetic planner.
@@
  • We are the first to explore the retrosynthetic planning task using an LLM alone. We introduce a comprehensive training pipeline integrating SFT and RL, enabling BioMedGPT-Mol to achieve state-of-the-art performance as an end-to-end retrosynthetic planner.
@@ 7 Further Exploring Retrosynthetic Planning with BioMedGPT-Mol
In contrast to these representation-focused approaches, our work investigates the potential of optimizing LLMs as standalone generative planners via explicit CoT reasoning. To the best of our knowledge, LLMs have not yet been successfully deployed as effective multi-step retrosynthetic planners in isolation. We demonstrate that BioMedGPT-Mol, empowered by its deep molecular expertise and optimized through SFT and RL, is capable of functioning as a standalone retrosynthetic planner, thereby bypassing the need for specialized external search algorithms.
In this section, we evaluate BioMedGPT-Mol as an end-to-end planner on the RetroBench dataset [liu2023fusionretro]. We adopt this benchmark for the rationality of its evaluation protocol, which employs a rigorous set-wise exact match metric. Unlike standard search success rates, this metric verifies whether the predicted starting materials can genuinely synthesize the target by checking for an exact match with the ground truth precursors. A prediction is deemed correct if it aligns with any feasible ground truth route. Furthermore, to ensure plausible route generation, the search is pruned if the predicted path length exceeds the depth of the ground truth pathway.
@@ 7.1 A Three-stage SFT Strategy for Retrosynthetic Planning
7.2 Reinforcement Learning with Verifiable Rewards
Following the supervised fine-tuning stages, we further optimize BioMedGPT-Mol using RL via Group Relative Policy Optimization (GRPO) [deepseekmath]. The primary objective of this stage is to enable the model to generalize beyond the rigid trajectories found in the training data. Retrosynthetic planning is inherently one-to-many; while SFT teaches the model to mimic specific annotated routes, RL encourages the exploration of solution spaces. Our goal is for the model to identify the correct set of commercially available starting materials and generate chemically plausible intermediate steps, even if the planned route differs from the specific ground-truth path provided in RetroBench.
To guide this optimization, we design a hierarchical reward function comprising three progressive components. Crucially, a subsequent reward component is evaluated and awarded only if the preceding constraint is fully satisfied.
  • Format Reward ($R_{\text{fmt}}$, weight=0.5): This foundational reward ensures the model adheres to the required output structure, specifically verifying the correct usage of <think> tags for reasoning and the interpretable format of the answer. If the format is incorrect, the evaluation terminates, and subsequent rewards are set to zero.
  • Validity Reward ($R_{\text{valid}}$, weight=0.5): Conditioned on a correct format, we verify the chemical validity of all generated molecular strings using RDKit. This reward penalizes the generation of invalid SMILES strings or chemically impossible structures.
  • Accuracy Reward ($R_{\text{acc}}$, weight=1.0): Conditioned on both correct format and valid chemical syntax, this metric calculates the set-wise exact match between the predicted starting materials and the ground truth precursors. A positive reward is granted if the predicted set aligns with the feasible solution defined in RetroBench.
The total reward $R$ is the accumulated sum of these weighted components, effectively guiding the model from basic structural coherence to chemical validity, and finally to planning accuracy.
@@ 7.3 Performance on RetroBench
Figure 4: Summary of retrosynthetic planning results for Top-1 exact match accuracy (%) on RetroBench.
A two-dimensional test-time scaling strategy is adopted, incorporating SMILES augmentation and weighted beam search. We generate $k_{a}$ distinct SMILES representations for the target molecule by varying the root atoms and traversal orders. For each augmented input, we subsequently perform a partial beam search with a width of $k_{b}$ to decode the most probable reactant SMILES. The final predictions are ranked by aggregating the local beam scores across all augmented inputs using the reciprocal rank sum:
$Score(y)=\sum_{j\in\mathcal{J}y}\frac{1}{r_{y,j}}$ (2)
where $y$ denotes a unique predicted reactant set, $\mathcal{J}_{y}$ comprises the indices of the augmented inputs where $y$ is generated, and $r_{y,j}$ represents the rank of candidate $y$ in the $j$-th partial beam search. In our experiments, we set $k_{a}=6$ and $k_{b}=2$.
Figure 4 illustrates the exact-match accuracy ($\%$) across various retrosynthetic planning methods. Our model, BioMedGPT-Mol, achieves a state-of-the-art accuracy of $43.6\%$, outperforming all comparative baselines. Notably, our method surpasses both the strongest semi-template-based/template-free competitor, RetroInText ($42.1\%$), and the leading template-based model, NeuralSym ($42\%$). This superior performance underscores the efficacy of our fully end-to-end, template-free architecture, demonstrating its capability to accurately predict retrosynthetic routes without relying on rigid reaction templates or external search algorithms.
@@ 8 Conclusion
In this work, we introduce BioMedGPT-Mol, a molecular language model designed for molecular understanding and generation. To create a molecule-centric research assistant and harness the power of open-source reasoning models with rapid iterative upgrades, we design and implement a multi-task learning strategy for model fine-tuning. Our experimental results demonstrate its remarkable performance, showing that a general-purpose reasoning model can be effectively and efficiently post-trained into a specialized molecular language model through a well-structured multi-task curriculum. Building on this foundation, we further explored the challenging task of retrosynthetic planning. By implementing a comprehensive training strategy that integrates SFT and RL, we enabled BioMedGPT-Mol to achieve state-of-the-art performance as an end-to-end retrosynthetic planner.